# Learning to Reconstruct Quirky Tracks

### v1

@@ IV Tracking
Most traditional algorithms make assumptions about the characteristics of the particles being reconstructed. This particle hypothesis typically includes the assumption that the particle track is approximately helical and that the particle has ionization and interaction properties typical of known particles. In this way, algorithms such as the kalman filter (KF) perform some form of helical fit as part of track finding. Machine learning approaches, such as graph neural networks (GNN), need not perform a fit and therefore need no such assumption. The characteristics of the particle tracks are learned by providing examples of target tracks, which may therefore be of any shape.
@@
There are a variety of ways ML can be used to solve the track finding problem (see Duarte and Vlimant (2022); DeZoort et al. (2023); Thais et al. (2022) for thorough reviews). It can be treated as an object detection problem and one can perform instance segmentation with object condensation Kieseler (2020); Lieret and DeZoort (2024); Murnane (2024). One can frame reconstruction as a generative process, where track candidates are recurrently produced by some autoregressive process Lad (2023); Huang et al. (2024). The combinatorial kalman filter (CKF) technique can also be combined with techniques for extending track seeds using ML Heinrich et al. (2024), or using reinforcement learning to choose a next hit Vaage (2022); Kortus et al. (2023).
@@
A successful approach has been to take advantage of a natural description of tracks as a set of directed graphs, and apply a GNN. The Exa.TrkX pipeline is built on the GNN paradigm, and uses the physics intuition of a track being a chain of sequentially connected hits Farrell et al. (2018); Ju et al. (2020b, 2021). This and similar approaches have been shown to give high track-finding efficiency and stable training on a diverse set of HEP experiments Caillou et al. (2022); Verma and Jena (2020); Hewes et al. (2021); Drielsma et al. (2021); Akram and Ju (2022); Jia et al. (2024); Biscarat et al. (2021); Zdybal et al. (2024); Andrews et al. (2021); DeZoort et al. (2021); Correia et al. (2023); Elabd et al. (2022). This pipeline contains three stages: a) Graph construction, where 3D spacepoints are connected into a graph of spacepoint-to-spacepoint edge connections; b) Edge classification, where edges are scored with a confidence of being either “true” (a connection between two sequential hits from a signal particle) or “fake” (any other kind of edge); and c) Graph segmentation, where low-scoring edges are removed, and collections of connected hits are used to produce track candidates.
@@
The Exa.TrkX pipeline can be applied nearly out-of-the-box to non-helical tracking, as in the case of quirk tracks. Some alterations to the definition of “true edges” generalize the pipeline to train correctly for this broader class of tracks. The most significant alteration is in the assumptions about the ordering of the hits. As many tracking detectors (as in ATLAS and TrackML) do not include reliable timing information, the ordering of hits in tracks within Exa.TrkX is defined heuristically using increasing distance from the production vertex of the track. This is reliable for high-momentum helical tracks that are unlikely to scatter back towards the interaction point. However, quirk tracks with a non-negligible opening angle are able to oscillate away from and towards the detector origin. Relying on the assumption that hits are in order according to their distance from the production vertex can lead to misinterpretation of the true track shape in training. In this work, we access the underlying simulation data to define the time-ordering of hits in order to correctly construct true edges for training, though such information is not required for testing.
@@
To construct graphs, we train a metric learning model, which attempts to embed sequential hits in a particle track into nearby positions in a latent space, while hits from other particles should be distant in the latent space. As this is purely data driven, tracks need not be helical, and any smooth particle trajectory may be learned by such a non-linear embedding. All that is required is a ground truth definition of pairwise similarity, as defined earlier. We use 150 training epochs and a learning rate of 0.001 in the metric learning for constructing the tracking graph. We then train a GNN to classify the edges as either “true” or “false”. In this step, we modify the epochs and learning rate for better efficiency and to avoid overfitting, using early stopping and decreasing the learning rate.
@@ V Track Reconstruction Performance
Performance of the tracking algorithm is measured in simulated samples described above. Performance is measured by tracking efficiency, the fraction of reconstructable particles that are reconstructed. Reconstructable particles are those at least 7 (22) true hits in the 8-layer (25-layer) geometry.
A particle is reconstructed if it is double-majority matched [^1] Double-majority matching requires that at least 50% of the reconstructed track hits belong to the truth track, and that at least 50% of the truth hits are found in the reconstructed track. to a matchable track candidate: one with least 5 (15) hits in the 8-layer (25-layer) geometry.
@@ V.1 Reconstructing quirk tracks using SM tracks
The SM-trained GNN-based tracker excels at finding SM tracks. For our benchmark sample, in the 8-layer geometry, SM tracks are found with 97.9% efficiency, while quirks are mostly missed, with 10.2% efficiency, using the definition of track efficiency above. For the 25 layer geometry, we find similar performance.
@@ V.2 Well-behaved quirks
Tracks of quirks are substantially more complex than helical tracks, and over the range of parameters can display a broad range of behavior that can be challenging for reconstruction algorithms. As seen in Fig. 1, tracks can cross individual layers multiple times due to oscillations. Before tackling these more difficult scenarios, we focus on the subset of quirk tracks identified as well-behaved by selecting those whose number of hits do not exceed the number of layers by a factor of three. This eliminates, for the moment, the most challenging quirk tracks, those which leave a very large number of hits as they travel in and out of detector layers, but retains tracks with clearly non-helical trajectories.
In the benchmark sample with 8 layers, training on a pure sample of well-behaved quirks produces a track finder with efficiency of 91.5% in events with only well-behaved quirks and no background tracks. This is a highly idealized scenario, but gives a rough upper bound on the performance of the pipeline and demonstrates its capacity to learn to find non-helical tracks when they comprise its training set, without background. Fig 3 shows an example of reconstructed and truth tracks in this scenario.
@@ V.3 All quirks
Quirks are an interesting and challenging problem precisely because they are not always well-behaved, and performance on these not-well-behaved tracks is also important for broad sensitivity. When the training sample is expanded to include all reconstructable tracks, the efficiency in our benchmark scenario with 8 detector layers is $56.3\%$ as compared to 7% for the tracker trained on SM tracks. Fig 4 presents an example of reconstructed and truth tracks.
@@ V.4 Finding Quirks among the Background
Realistic applications of non-helical tracking require identifying such tracks among many background helical tracks from SM processes. To assess the performance in this environment, we use a training sample which includes both quirks with positive labels as well as SM tracks with negative labels. The efficiency in this scenario is $71.9$% for all quirk tracks in our benchmark sample with eight detector layers, somewhat less efficient than the much simpler task of background-free tracking, which scores 91.5%.
@@ V.5 Dependence on Geometry
To explore the dependence of the number of tracking layers, the study was repeated in the 25-layer setting, yielding an efficiency of $79.0\%$, somewhat higher than the efficiency in the 8-layer case. This suggests that the nature of the problem does not sensitively depend on the number of layers. Figure 6 presents an example of reconstructed and truth tracks with 25 layers for $m_{Q}=500$ GeV, $\Lambda=500$ eV.
Figure 6: Example truth and reconstructed quirk tracks for $m_{Q}=500$ GeV, $\Lambda=500$ eV in the 25 layer geometry.
@@ V.6 Dependence on Quirk Parameters
For very long oscillation lengths, ${\cal O}({\rm m})$, no oscillations occur in the tracking volume and the problem is essentially indistinguishable from that of the SM tracking, and efficiencies are very high, $\approx 90$%. For the target regime of intermediate-length oscillation lengths, ${\cal O}({\rm cm})$, performance is in the range of $60-80$%, which is a dramatic improvement over the $\approx 10\%$ efficiency for quirk tracks when using a SM-trained tracker, see Tab. 2. As the oscillation length becomes very small, reported efficiency drops by a factor of $\approx 2$; this is due to the two quirk tracks overlapping nearly exactly, often leading to only one reconstructed track, though in principle two exist; see Fig. 7.
@@ V.7 Generalization across Quirk Parameters
The efficiency to reconstruct quirks generated with $m_{Q}=100$ GeV, $\Lambda=100$ eV is 52.6% when training with a mixture that does not include that set, compared to 77.8% when training only on tracks from the same parameter set as the testing sample. This demonstrates some significant but imperfect ability to generalize. Another working point with $m_{Q}=1000$ GeV, $\Lambda=1000$ eV achieves 49.3% efficiency for the generalized training, compared to 62.7% for the specific training, a similar performance ratio. The nature of the tracking challenge varies with quirk parameter values, but each working point clearly has enough in common for the tracker to learn general features of quirk tracks, not just the features of quirks for one particular quirk parameter choice. Further work is required to understand which ML choices contribute to the highest generalizability.
Figure 7: Examples of truth and reconstructed tracks for well-behaved quirks with $m_{Q}=500$ GeV, $\Lambda=5000$ eV, in which there is no visible oscillation and the two quirks overlap.
V.8 Hit Location Uncertainty and Pileup Impact
To explore the sensitivity of the pipeline to imperfect detector resolution, we smear the hit locations with a Gaussian of width 1mm, a reasonable approximation of LHC tracking resolution Mindur (2017). A spot test for $m_{Q}=500$ GeV, $\Lambda=500$ eV quirks shows no significant degradation in tracking performance, with performance consistent within statistical uncertainties.
To explore the sensitivity of the pipeline to the expected LHC environment, which includes many $pp$ interactions, we include minimum bias events generated with Pythia. The inclusion of 5-10 pileup events degrades performance by approximately 5-10%.
@@ V.8 Hit Location Uncertainty and Pileup Impact
VI Discussion
@@
Reconstruction of quirk tracks in the region where the oscillation length ($d$) is larger than detector resolution, but smaller than the detector is expected to be very challenging. The application of ML-based tracking demonstrates a new-found capability to reconstruct such tracks with reasonable efficiency, in excess of 50% across a range of parameter values where $d\approx 1-10$ cm. For comparison, previous work which searches for co-planar hits to avoid fitting quirk trajectories Knapen et al. (2017) has a similar reconstruction efficiency for small oscillation lengths, $d<1$ cm, but drops very quickly for cm scale oscillations, dipping down to $\approx 10\%$ for $m_{Q},\Lambda$ = 800 GeV, 1000 eV. Training ML trackers explicitly on quirks can significantly extend the sensitivity range, complementing the co-planar approach.
@@
One approach to implementation of an ML-based quirk-finding algorithm is to analyze a set of events which are selected for other reasons, such as large $p_{\rm T}^{\rm miss}$. But a more ambitious and much more challenging direction would be to generalize the core track-finding in the reconstruction pipeline so that it is capable of identifying helical and non-helical tracks, whether those are quirks or other unexpected particles.
@@
Ref. Knapen et al. (2017) estimated the trigger efficiency for quirks directly to be low with existing tools, relying on primarily muon triggers. However, an ML-based tracker for quirks opens a new possibility of identifying quirks directly in the trigger, using knowledge distillation and hardware accelerators to fit the GNN within the trigger budget.
VII Conclusions

### v2

@@ IV Tracking
Most traditional algorithms make assumptions about the characteristics of the particles being reconstructed. This particle hypothesis typically includes the assumption that the particle track is approximately helical and that the particle has ionization and interaction properties typical of known particles. In this way, algorithms used in collider tracking such as the kalman filter (KF) propagate a helical (or otherwise modeled) state vector and update it hit-by-hit as part of track finding. Machine learning approaches, such as graph neural networks, need not explicitly model this state and therefore need no such assumption. The characteristics of the particle tracks are learned by providing examples of target tracks, which may therefore be of any shape.
@@
There are a variety of ways ML can be used to solve the track finding problem (see Duarte and Vlimant (2022); DeZoort et al. (2023); Thais et al. (2022) for thorough reviews). It can be treated as an object detection problem and one can perform instance segmentation with object condensation Kieseler (2020); Lieret and DeZoort (2024); Murnane (2024). One can frame reconstruction as a generative process, where track candidates are recurrently produced by some autoregressive process Huang et al. (2024). The combinatorial kalman filter (CKF) technique can also be combined with techniques for extending track seeds using ML Heinrich et al. (2024), or using reinforcement learning to choose a next hit Vaage (2022); Kortus et al. (2023).
@@
A successful approach has been to take advantage of a natural description of tracks as a set of paths in a directed graph, and apply a GNN. The Exa.TrkX pipeline is built on this GNN paradigm, and uses the physics intuition of a track being a chain of sequentially connected hits Farrell et al. (2018); Ju et al. (2020b, 2021). This and similar approaches have been shown to give high track-finding efficiency and stable training on a diverse set of HEP experiments Caillou et al. (2022); Verma and Jena (2020); Hewes et al. (2021); Drielsma et al. (2021); Akram and Ju (2022); Jia et al. (2024); Biscarat et al. (2021); Zdybal et al. (2024); Andrews et al. (2021); DeZoort et al. (2021); Correia et al. (2024); Elabd et al. (2022). This pipeline contains three stages: a) Graph construction, where 3D spacepoints are connected into a graph of spacepoint-to-spacepoint edge connections; b) Edge classification, where edges are scored with a confidence of being either “true” (a connection between two sequential hits from a signal particle) or “fake” (any other kind of edge); and c) Graph segmentation, where low-scoring edges are removed, and collections of connected hits are used to produce track candidates.
@@
The Exa.TrkX pipeline can be applied nearly out-of-the-box to non-helical tracking, as in the case of quirk tracks. Some alterations to the definition of “true edges” generalize the pipeline to train correctly for this broader class of tracks. The most significant alteration is in the assumptions about the ordering of the hits. As many tracking detectors (as in ATLAS and the open detector used in the TrackML challenge Kiehn et al. (2019); Amrouche et al. (2019, 2021)) do not include reliable timing information, the ordering of hits in tracks within Exa.TrkX is defined heuristically using increasing distance from the production vertex of the track. This is reliable for high-momentum helical tracks that are unlikely to scatter back towards the interaction point. However, quirk tracks with a non-negligible opening angle are able to oscillate away from and towards the detector origin. Relying on the assumption that hits are in order according to their distance from the production vertex can lead to misinterpretation of the true track shape in training. In this work, we access the underlying simulation data to define the time-ordering of hits in order to correctly construct true edges for training, though such information is not required for testing.
@@
To construct graphs, we train a metric learning model, which attempts to embed sequential hits in a particle track into nearby positions in a latent space, while hits from other particles should be distant in the latent space. As this is purely data driven, tracks need not be helical, and any smooth particle trajectory may be learned by such a non-linear embedding. All that is required is a definition of pairwise similarity, as defined earlier. In our training analyses, we split our dataset into training, validation, and testing sets with 2k, 400, and 400 samples respectively. We use 150 training epochs and a learning rate of 0.001 in the metric learning for constructing the tracking graph. We then train a GNN to classify the edges as either “true” or “false”. In this step, we modify the epochs and learning rate for better efficiency and to avoid overfitting by checking the loss and efficiency plots, using early stopping and decreasing the learning rate. Also, we define the “score cut” to select our target candidates. This value is set to 0.9 in most analyses and is optimized in some cases.
@@ V Track Reconstruction Performance
Performance of the tracking algorithm is measured in simulated samples described above. Performance is measured by tracking efficiency, the fraction of reconstructable particles that are reconstructed. Reconstructable particles are those with at least 7 (22) true hits in the 8-layer (25-layer) geometry. A particle is reconstructed if it is double-majority matched [^1] Double-majority matching requires that at least 50% of the reconstructed track hits belong to a matching truth track, and that at least 50% of the hits in that truth track are found in the reconstructed track. to a matchable track candidate: one with least 5 (15) hits in the 8-layer (25-layer) geometry.
@@ V.1 Reconstructing quirk tracks using SM tracks
The SM-trained GNN-based tracker excels at finding SM tracks. For our benchmark sample, in the 8-layer geometry, SM tracks are found with 97.9% efficiency, while quirks are mostly missed, with 10.2% efficiency, using the definition of track efficiency above. For the 25-layer geometry, we find similar performance.
@@ V.2 Well-behaved quirks
Tracks of quirks are substantially more complex than helical tracks, and over the range of parameters can display a broad range of behavior that can be challenging for reconstruction algorithms. As seen in Fig. 1, tracks can cross individual layers multiple times due to oscillations. Before tackling these more difficult scenarios, we focus on the subset of quirk tracks identified as well-behaved, those whose number of hits do not exceed the number of layers by a factor of three. This eliminates, for the moment, the most challenging quirk tracks, those which leave a very large number of hits as they travel in and out of detector layers, but retains tracks with clearly non-helical trajectories.
In the benchmark sample with 8 layers, training on a pure sample of well-behaved quirks produces a track finder with efficiency of 91.5% in events with only well-behaved quirks and no background tracks. This is a highly idealized scenario, but gives a rough upper bound on the performance of the pipeline and demonstrates its capacity to learn to find non-helical tracks when they comprise its training set, without background. Figure 3 shows an example of reconstructed and truth tracks in this scenario.
@@ V.3 All quirks
Quirks are an interesting and challenging problem precisely because they are not always well-behaved, and performance on these not-well-behaved tracks is also important for broad sensitivity. When the training sample is expanded to include all reconstructable tracks, the efficiency in our benchmark scenario with 8 detector layers is $56.3\%$ as compared to 10.2% for the tracker trained on SM tracks. Figure 4 presents an example of reconstructed and truth tracks.
@@ V.4 Finding Quirks among the Background
Realistic applications of non-helical tracking require identifying such tracks among many background helical tracks from SM processes. To assess the performance in this environment, we use a training sample which includes tracker hits from both quirk and SM particles in each event. In graph construction and edge classification, only quirk hit pairs are treated as true. All other pairs, including pairs of hits from the same SM particle, are treated as fake in training, and models are encouraged to not reconstruct them. Previous work has explored other strategies with such background tracks, such as treating them as true or masking them from the loss function Caillou et al. (2022). In this work, our goal is quirk finding at the highest possible background rejection rate, rather than the reconstruction of all true tracks, and this motivates the strategy as described. The efficiency in this scenario is 71.971.9% for all quirk tracks in our benchmark sample with eight detector layers, somewhat less efficient than the much simpler task of background-free tracking, which scores 91.5%. Background tracks are also included in training discussed in subsequent studies in the following sections, as well as the results in Tab. 1.
@@ V.5 Dependence on Geometry
To explore the dependence of the number of tracking layers, the study was repeated in the 25-layer setting, yielding an efficiency of $79.0\%$, somewhat higher than the efficiency in the 8-layer case. This suggests that the nature of the problem does not sensitively depend on the number of layers. Figure 6 presents an example of reconstructed and truth tracks with 25 layers for $m_{Q}=500$ GeV, $\Lambda=500$ eV. The 25-layer geometry is used in studies in the following sections.
Figure 6: Example truth and reconstructed quirk tracks for $m_{Q}=500$ GeV, $\Lambda=500$ eV in the 25 layer geometry. The ”X” represents truth hits.
@@ V.6 Dependence on Quirk Parameters
For very long oscillation lengths, ${\cal O}({\rm m})$, no oscillations occur in the tracking volume and the problem is essentially indistinguishable from that of the SM tracking, and efficiencies are very high, $\approx 90$%. For the target regime of intermediate-length oscillation lengths, ${\cal O}({\rm cm})$, performance is in the range of 608060-80%, which is a dramatic improvement over the $\approx 10\%$ efficiency for quirk tracks when using a SM-trained tracker (compare Tab. 1 Tab. 2). As the oscillation length becomes very small, reported efficiency drops by a factor of $\approx 2$; this is due to the two quirk tracks overlapping nearly exactly, often leading to only one reconstructed track, though in principle two exist; see Fig. 7.
@@ V.7 Generalization across Quirk Parameters
The efficiency to reconstruct quirks generated with $m_{Q}=100$ GeV, $\Lambda=1000$ eV is 52.6% when training with a mixture that does not include that set, compared to 77.8% when training only on tracks from the same parameter set as the testing sample. The mixed set includes $m_{Q}~({\rm GeV}),\Lambda~({\rm eV})=(100,100),(500,100),(1000,100)(100,500)(5000,1000)$. This demonstrates some significant but imperfect ability to generalize. Another working point with $m_{Q}=1000$ GeV, $\Lambda=1000$ eV achieves 49.3% efficiency for the generalized training, compared to 62.7% for the specific training, a similar performance ratio. The nature of the tracking challenge varies with quirk parameter values, but each working point clearly has enough in common for the tracker to learn general features of quirk tracks, not just the features of quirks for one particular quirk parameter choice. Further work is required to understand which ML choices contribute to the highest generalizability.
Figure 7: Examples of truth and reconstructed tracks for well-behaved quirks with $m_{Q}=500$ GeV, $\Lambda=5000$ eV, in which there is no visible oscillation and the two quirks overlap. The ”X” represents truth hits.
V.8 Hit Location Uncertainty
To explore the sensitivity of the pipeline to imperfect detector resolution, we smear the hit locations in training and testing with a Gaussian of width 1 mm, which is larger than typical LHC tracking resolution Mindur (2017); Chatrchyan et al. (2014). A spot test for $m_{Q}=500$ GeV, $\Lambda=500$ eV quirks shows no significant degradation in tracking performance, with performance consistent within statistical uncertainties.
@@ V.8 Hit Location Uncertainty
VI Backgrounds and Sensitivity
The pipeline has demonstrated the capacity to find quirks with a reasonable efficiency, which is necessary to observe quirks in experiments at the LHC. However, equally important to control is the background (or ”fake”) rate. Quirks are an exciting target because their signature is striking enough to suggest the possibility of very low background, potentially even background-free, discoveries. Unfortunately, SM tracks have a very high rate, so even a tiny probability to mimic the quirk signature could significantly impact discovery potential.
To examine the rate of background from SM tracks, we collect samples of fake tracks reconstructed by the pipeline which are not matched to quirks. We perform two fits to each track, one to the quirk hypothesis and a second to the helix hypothesis and use the ratio of $\chi^{2}$ values to discriminate between quirks and the background. In both the quirk and helix hypothesis, the $\chi^{2}$ is the sum of the squared distances between a hit and the distance of closest approach by the track, with an implicit $\sigma^{2}=1$ cm2 for simplicity.
The helix fitter finds the value of the five helical parameters that minimize the $\chi^{2}_{\textrm{helix}}$, starting from an initial estimate Hansroul et al. (1988) and refined by the Nelder-Mead optimization algorithm. This approach does not match the sophistication of track fitting in a realistic setting, which benefits from additional heuristics, such as outlier removal and ambiguity resolution. For our purposes, this provides a conservative estimate on our ability to distinguish between true quirks and background. The distribution of $\chi^{2}_{\textrm{helix}}$ is shown in the upper left panel of Fig. 8 for true quirks at our benchmark point $(m_{Q}=500,\Lambda=500)$ and for background particles. There is a clear separation between the two distributions, as many quirks cannot be well fit to a helix. A significant population of background tracks also receive a $\chi^{2}_{\textrm{helix}}$ larger than what might be expected, due to incorrect hits and local minima in the fitting.
In the quirk hypothesis, we use the propagation algorithm as described in Sec.III, which requires that we fit simultaneously for the four-vectors of the quirk pair given values of $(m_{Q},\Lambda)$, though only one track is required and included in the $\chi^{2}$ evaluation. The distribution of the ratio of $\chi^{2}_{\textrm{quirk}}$ to $\chi^{2}_{\textrm{helix}}$ values are shown in the left panel of Fig. 8 for true quirks and for background SM tracks. Both background tracks can be accomodated by the quirk hypothesis, due to the freedom provided by the second quirk’s unconstrained parameters. To mimic a high-momentum SM track, the fitter finds solutions in which the two quirks are back to back, generating perfectly straight line. To suppress these, we remove tracks whose fitted quirks are back to back ($\Delta\phi>2.8$) or whose ratio of $\ln({\chi^{2}_{\textrm{quirk}}/\chi^{2}_{\textrm{helix}}})<-1$; see the lower panel of Fig. 8. For $m_{Q}=500$ GeV, $\Lambda=500$ eV, this requirement is 94% efficient for quirks and removes all SM tracks in our samples. For other $m_{Q},\Lambda$ comparable efficiencies of $\approx 80\%$ are found, except for the heaviest case $m_{Q}=5000$ GeV, due to their larger opening angles (however, 5 TeV quirks have a small cross section and are not produced at any appreciable rate). To ensure a negligible background, these requirements can be tuned in data, but these studies suggest that such suppression is possible. Further suppression is available by fitting two quirk candidate tracks simultaneously, if needed. Additional discrimination power may be available from the momentum of the recoiling jet, which is correlated with the quirk pair momentum.
To estimate the potential of discovering the quirk signal among the SM background, we estimate the number of reconstructed quirk events, $N_{\rm Reconstructed}$, as follows. Following Ref. Knapen et al. (2017), we select quirk events that include a $p_{\textrm{T}}>200$ GeV jet to pass the High Level Trigger, and calculate the number of reconstructed quirk events as
$N_{\rm Reconstructed}={\cal L}\times\sigma(Q\bar{Q}j,p_{\textrm{T}}^{\textrm{jet}}>200~{\rm GeV})\times\epsilon_{\rm total},$ (3)
where $\epsilon_{\rm total}$ is the total efficiency for reconstruction, selection and also includes the fraction of well-behaved quirks at this parameter point. The efficiency $\epsilon_{\rm total}$ varies from $\lesssim$1% to $\approx$ 30% across the plane as shown in the left panel of Fig. 9. The right panel of Fig 9 shows the expected quirk event rate $N_{\rm Reconstructed}$ in 300 fb-1 of $pp$ collisions at the LHC, showing $N>3$ across $\Lambda\lesssim 5000$ eV for $m_{Q}\lesssim 700$ GeV.
Figure 8: Distributions of ratio $\chi^{2}$ under the quirk and helix hypothesis, for $m_{Q}=500~{\rm GeV},\Lambda=500~{\rm eV}$quirk and SM tracks. The lower panel shows the ratio of the two $\chi^{2}$ metrics versus the fitted quirk opening angle for the same $m_{Q},\Lambda$, and the selection used to remove SM tracks is shown in the dashed red lines.
Figure 9: Left: Efficiency of quirk finding versus mass $m_{Q}$ and string tension $\Lambda$. Right: Expected reconstructed quirk event rate per 300 fb-1 in the same parameter plane.
VII Discussion
@@
Reconstruction of quirk tracks in the region where the oscillation length ($d$) is larger than detector resolution, but smaller than the detector is expected to be very challenging. The application of ML-based tracking demonstrates a new-found capability to reconstruct such tracks with reasonable efficiency, in excess of 50% across a range of parameter values where $d\approx 1-10$ cm. For comparison, previous work which searches for co-planar hits to avoid fitting quirk trajectories Knapen et al. (2017) has a similar reconstruction efficiency for small oscillation lengths, $d<1$ cm, but drops very quickly for cm scale oscillations, reaching $\approx 10\%$ for $m_{Q},\Lambda$ = 800 GeV, 1000 eV. Training ML trackers explicitly on quirks can significantly extend the sensitivity range, complementing the co-planar approach. The technique presented in Sec. V is found to be sensitive to $m_{Q}\lesssim 700$ GeV, for $100{~\rm eV}<\Lambda<5000{~\rm eV}$.
@@
One approach for the implementation of an ML-based quirk-finding algorithm is to analyze a set of events which are selected for other reasons, such as large $p_{\rm T}^{\rm miss}$. But a more ambitious and much more challenging direction would be to generalize the core track-finding in the reconstruction pipeline so that it is capable of identifying helical and non-helical tracks, whether those are quirks or other unexpected particles.
@@
Ref. Knapen et al. (2017) estimated the trigger efficiency for quirks directly to be low with existing tools, relying on primarily muon triggers. However, an ML-based tracker for quirks opens a new possibility of identifying quirks directly in the trigger, using knowledge distillation  Hinton et al. (2015) and hardware accelerators such as GPUs to fit the GNN within the trigger budget.
VIII Conclusions